OpenStudy (anonymous):

Solve the inequality. Show your work. 6y-8<10

OpenStudy (anonymous):

solve it like you would solve an equation just remember for inequalities, if you ever multiply or divide a negative number to both sides, you must change the sign

OpenStudy (anonymous):

\[6y-8\le10\]

OpenStudy (anonymous):

pretend that it is \[6y-8 = 10\] and solve it

OpenStudy (anonymous):

what do you think your first step would be?

OpenStudy (anonymous):

please understand this is my first time taking Algebra

OpenStudy (anonymous):

... why are you doing inequalities when you do not know how to solve equations....

OpenStudy (anonymous):

ex. solving an equation 8x-6=2 the goal is to isolate the variable so its alone on one side the first step would be to remove the -6 from the left side this can be accomplished by adding 6 to both sides by doing this 8x-6 +6 = 2+6 -6+6 = 0 so the -6 cancels out on the left 8x = 8 now at this point, we want to isolate x so it alone on one side this means we need to get rid of the 8 on the left side this can be accomplished with we divide 8 to both sides because 8/8 = 1 and 1x is the equivalent of just x so the equation becomes x= 8/8 and 8/8 = 1 now see if you can apply these steps to your own problem

OpenStudy (anonymous):

one more note i mentioned that when dividng by a negative number to both sides of an inequality causes the sign to change this is what i meant \[-x< 1\] dividing by -1 to both sides or multiplying \[x> -1\]